# Joint Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems: A DRL Approach

### v1

@@ IV-D Latency Performance Analysis
Figure 10: Average delay performance as distinct user’s arrival rate $\lambda$ varies.
IV-E Impact of Reflection Elements $M$
@@
IV-F Latency Analysis in Brustiness Scenarios
To evaluate the performance of the proposed algorithm in burstiness scenarios, we introduce sudden data arrivals at specific time steps. To simulate this situation, we consider that buffers 1, 2, and 3 receive 300 data packets at time slots 100, 400, and 700, respectively, to emulate bursty traffic. Fig. 8 presents a comparison of the delay performance between the proposed algorithm and the “max minimum rate” algorithm under different scenarios. It can be observed that, regardless of whether burstiness is present, the proposed algorithm consistently demonstrates superior performance, effectively validating its effectiveness, in handling burstiness. To illustrate why the proposed algorithm exhibits superior performance in burstiness scenarios, a specific case is selected, and the number of backlogged packets in buffers and transmission rates are analyzed. As shown in Fig. 9, Fig. 9LABEL:sub@Fig:packagePropsed and Fig. 9LABEL:sub@Fig:packageMaxMin depict the backlogged packets in each buffer, while Fig. 9LABEL:sub@rateProposed and Fig. 9LABEL:sub@rateMaxMin present the corresponding user rates, where the first 300 steps are extracted.
As illustrated in Fig. 9LABEL:sub@Fig:packagePropsed, the proposed algorithm effectively reduces the number of backlogged packets caused by burstiness through dynamic adjustment of the resource allocation strategy. In contrast, Fig. 9LABEL:sub@Fig:packageMaxMin shows that under the “max minimum rate” algorithm, the number of backlogged packets in the buffer 3 remains high for an extended period, indicating inefficient resource allocation strategy. Furthermore, Fig. 9LABEL:sub@rateProposed demonstrates that the proposed algorithm allocates more resources to users with a larger packet backlog, leading to an immediate increase in their transmission rates. This is particularly evident for the user 3, whose rate surges, allowing the backlogged packets to be quickly transmitted, thereby reducing latency. In contrast, Fig. 9LABEL:sub@rateMaxMin reveals that the “max minimum rate” method distributes resources more rigidly, ingoring the status of buffers and preventing heavily congested users from transmitting their backlog efficiently, resulting in prolonged delays.
@@ IV-G Latency Analysis under Different Traffic Flows
To evaluate the performance of the proposed algorithm under different user arrival rates. We consider the packet arrival rates $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ to be $\lambda+\text{gap}$, $\lambda$, and $\lambda-\text{gap}$, respectively. As shown in Fig. 10, the delay of the “max minimum rate” method increases as the gap grows, while the proposed algorithm maintains a relatively low delay across different gap values. It implies that, although the proposed algorithm was trained when the gap was 0, it exibits strong robustness in different traffic scenarios. To evaluate the superiority of the proposed algorithm in handling different traffic flows, we consider a representative case where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 12, 9.5, and 7, respectively. To provide a clear illustration, we focus on the first 300 time slots, with an emergency burst of 300 packets arriving for the buffer 3 at the time slot 100. As shown in Fig. 11, Fig. 11LABEL:sub@Fig:packagePropsed and Fig. 11LABEL:sub@Fig:packageMaxMin display the number of backlogged packets in each user’s buffer, while Fig. 11LABEL:sub@rateProposed and Fig. 11LABEL:sub@rateMaxMin illustrate the corresponding transmission rates.
From Fig. 11LABEL:sub@Fig:packagePropsed, it can be observed that despite different arrival rates among users, the numbers of backlogged packets in the buffers are approximately balanced overall, demonstrating the fairness of the proposed algorithm. At the time slot 100, when buffer 3 experiences a sudden arrival of packets, the algorithm effectively responds by dynamically adjusting resource allocation, allowing the accumulated packets in the buffer 3 to be transmitted promptly. In contrast, as shown in Fig. 11LABEL:sub@Fig:packageMaxMin, the “max minimum rate” algorithm continues to allocate resources based solely on rate fairness, which results in severe buffer congestion for high arrival rate users such as User 1. Further insights can be drawn from Fig. 11LABEL:sub@rateProposed and Fig. 11LABEL:sub@rateMaxMin, where the advantage of the proposed algorithm becomes more evident. It adaptively allocates resources based on users’ arrival rates, assigning fewer resources to low arrival rate users and more resources to high arrival rate users. Additionally, the algorithm dynamically balances the buffer congestion and arrival rates to optimize resource allocation and minimize delay. This is particularly noticeable before the time slot 100 in Fig. 11LABEL:sub@rateProposed, where the user 3 initially has a lower transmission rate due to its relatively low arrival rate. However, when a burst occurs, the system promptly allocates more resources, enabling the user 3 to efficiently clear the backlog in its buffer. In contrast, the “max minimum rate” method, as illustrated in Fig. 11LABEL:sub@rateMaxMin, strictly enforces rate fairness without considering the number of backlogged packets and varying arrival rates across users. This leads to inefficient resource allocation, as users with excessive backlog receive insufficient resources, resulting in prolonged delays. Consequently, the proposed algorithm outperforms the “max minimum rate” algorithm by providing adaptive and delay-aware resource allocation, ensuring both fairness and efficiency in dynamic traffic scenarios.

### v3

@@ IV-D Latency Performance Analysis
Figure 10: Impact of reflection elements $M$ on delay performance.
IV-E Latency Analysis in Brustiness Scenarios
To evaluate the proposed algorithm under bursty traffic, we simulate sudden data arrivals by injecting 300 packets into buffers 1, 2, and 3 at time slots 100, 400, and 700, respectively. Fig. 7 presents a comparison of the delay performance between the proposed algorithm and the “max minimum rate” algorithm under different scenarios. It can be observed that, regardless of whether burstiness is present, the proposed algorithm consistently demonstrates superior performance, effectively validating its effectiveness, in handling burstiness.
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. As shown in Fig. 8LABEL:sub@Fig:packagePropsed, the proposed algorithm effectively reduces backlog by dynamically adjusting resource allocation in response to bursty arrivals. In contrast, Fig. 8LABEL:sub@Fig:packageMaxMin shows persistent congestion in buffer 3 under the “max minimum rate” approach, indicating inefficient allocation. Fig. 8LABEL:sub@rateProposed further shows that the proposed method prioritizes users with heavier backlogs, rapidly increasing their transmission rates, such as user 3, leading to quick backlog clearance. Conversely, Fig. 8LABEL:sub@rateMaxMin reveals that the baseline allocates resources more rigidly, ignoring buffer states and failing to relieve congestion effectively, which results in higher delay.
IV-F Impact of Reflection Elements $M$
@@ IV-G Latency Analysis under Different Traffic Flows
Figure 11: Average delay performance as distinct packet arrival rate $\lambda$ varies.
To evaluate the performance of the proposed algorithm under varying user arrival rates, we set the packet arrival rates $\lambda_{1}=\lambda+\text{gap}$, $\lambda_{2}=\lambda$, and $\lambda_{3}=\lambda-\text{gap}$. As shown in Fig. 11, the delay of the “max minimum rate” method increases with larger gaps, while the proposed algorithm maintains consistently low delay, indicating strong robustness even when trained under uniform arrival rates.
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. In Fig. 9LABEL:sub@Fig:packagePropsed, the proposed algorithm keeps buffer backlogs balanced despite differing arrival rates, showing its fairness. Upon the burst at time slot 100, it quickly reallocates resources to clear buffer 3. In contrast, Fig. 9LABEL:sub@Fig:packageMaxMin shows the “max minimum rate” approach leads to severe congestion in buffer 1 due to static rate-based allocation. Further, Fig. 9LABEL:sub@rateProposed reveals the proposed method allocates more resources to high arrival rate users and dynamically adapts to queue states. Before time slot 100, user 3 has a lower rate due to low arrival, but receives rapid support when the burst occurs. Conversely, Fig. 9LABEL:sub@rateMaxMin shows that the “max minimum rate” method rigidly enforces rate fairness, ignoring buffer states, resulting in inefficient resource use and longer delays.